OpenStudy (anonymous):

3 2/3 = what square root

OpenStudy (anonymous):

question really is not very clear is that exactly what it says?

OpenStudy (anonymous):

Ok what is 3 2/3 equal to? 3 square root 9 not sure if I'm writing it correctly 2 square root 9 3 square root 9

OpenStudy (anonymous):

it is really \[\huge 3^{\frac{2}{3}}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok the exponent is \(\frac{2}{3}\) the numerator is \(2\) that means square it the denominator is \(3\) , that means take the cubed root

OpenStudy (anonymous):

so \[\huge 3^{\frac{2}{3}}=\sqrt[3]{3^2}\]

OpenStudy (anonymous):

or if you prefer \[\sqrt[3]{9}\]

OpenStudy (anonymous):

Ok thank you very much.

OpenStudy (anonymous):

Ok have another question which is division Simplify 3 square root 7 divided by 5 square root 7 7 1/5 7 5/3 7 2/15